Two events E1 and E2 associated with a sample space of an experiment are such that P(E1)=0.27, P(E2)=0.40 and P(E1 U E2)=0.58. Determine whether E1 and E2 are independent.


1
Expert's answer
2021-07-19T05:46:31-0400
"P(E_1\cup E_2)=P(E_1)+P(E_2)-P(E_1\cap E_2)"

"P(E_1\cap E_2)=P(E_1)+P(E_2)-P(E_1\cup E_2)"

"=0.27+0.40-0.58=0.09"

"P(E_1)P(E_2)=0.27(0.40)=0.108"

"\not=0.09=P(E_1\cap E_2)"

Since "P(E_1\cap E_2)\not=P(E_1)P(E_2)," then "E_1" and "E_2" are not independent.